OpenStudy (anonymous):

I'm not too sure how to approach this question. Use Power series operations to find the Taylor series at x = 0 for the given function. f(x) = x^2e^x

OpenStudy (anonymous):

\[f(x) = x^2e^x\]

OpenStudy (anonymous):

f(x) = f(0) + f'(x)(x-0) + (f''(x)(x-0)^2)/2! +(f'''(x)(x-0)^3)/3! + ...

OpenStudy (mathmale):

Here you have a function which has the form of a product. Please write out the power series for the exponential function, y=e^x. Once you have a formula for that, you need only multiply it (or multiply each term) by x^2. pgpilot is offering you another approach which is correct; it's the general form of a Taylor series with center at x=0.

OpenStudy (anonymous):

oh, e^x = 1+x + x^2/2! + ... just multiply out

OpenStudy (anonymous):

Great. thanks for the help